# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ B.1. Conjugated operator
For all $i\in\llbracket 1,n\rrbracket$, let us consider the functions $\gamma_{i}$ such that $\mbox{reg}(\gamma)<\mbox{reg}^{0}$ and the following notation
$\nabla_{\gamma}f\triangleq(\sqrt{\gamma_{1}}D_{1}f_{1},\cdots,\sqrt{\gamma_{n}}D_{n}f_{n})\quad\text{and}\,\quad\Delta_{\gamma}f\triangleq\sum_{i=1}^{n}\gamma_{i}\partial_{x_{i}}^{2}f.$
Let $\mathcal{P}(w)\triangleq\,dw-\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}w\right)dt=fdt+gdB(t)$. With respect to this weight function, the first step is to consider the change of variable $w=\rho z$. Our first task is to split the conjugate operator
@@
into simple terms that we will estimate separately. Using Ito’s formula (cf. (klebaner2012introduction, , Theorem 4.18)) and notice that $r\rho=1$, we obtain the following result
@@
(B.1) $\begin{split}rd(\rho z)-r\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}(\rho z)\right)dt+R_{h}z\,dt=&B_{1}z+B_{2}z\,dt+C_{1}z\,dt+C_{2}z\,dt+C_{3}z\,dt,\end{split}$
@@
$\begin{split}R_{h}z\triangleq\,&\sum_{i=1}^{n}\left(h\mathcal{O}(1)rD_{i}^{2}\rho+D_{i}\gamma_{i}rA_{i}D_{i}\rho\right)A_{i}^{2}z+\sum_{i=1}^{n}h_{i}\mathcal{O}(1)rD_{i}A_{i}\rho D_{i}A_{i}z\\ &+\sum_{i=1}^{n}\frac{h^{2}}{4}D_{i}\gamma_{i}rD_{i}^{2}\rho D_{i}A_{i}z+\sum_{i=1}^{n}\frac{h^{2}}{4}D_{i}\gamma_{i}\,rD_{i}A_{i}\rho D_{i}^{2}z.\end{split}$
@@
$C_{4}z\triangleq\,-\frac{h^{2}}{4}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(rD_{i}^{2}\rho)A_{i}z),\quad C_{5}z\triangleq\,-\frac{h^{2}}{4}\sum_{i=1}^{n}D_{i}(D_{i}(\gamma_{i}rD^{2}_{i}\rho)A_{i}z)$
$C_{6}z=-B_{3}z\triangleq\,-2s\Delta_{\gamma}\varphi\,z$
in (B.1) we have the following identity in $Q$
(B.2) $r\mathcal{P}(\rho z)=Cz\,dt+Bz-M_{h}z\,dt,$
@@
Thus, our next task is to estimate the cross-product $2\mathbb{E}\int_{Q}CzBz=2\mathbb{E}\sum_{i=1}^{6}\sum_{j=1}^{3}\int_{Q}C_{i}zB_{j}z$ and to bound the left-hand side of (B.2) with respect to the term $M_{h}z$. Indeed, We notice that from (B.2), it holds that
(B.3) $2r\mathcal{P}(q)\,Cz=2|Cz|^{2}dt+2Cz\,Bz-2M_{h}z\,Cz\,dt.$
Hence, combining the definition of $r\mathcal{P}(q)$ with (B.3), it follows that
@@
(B.4) $2\mathbb{E}\int_{0}^{T}|rf|^{2}\,dt\geq\mathbb{E}\int_{Q}|Cz|^{2}dt+2\mathbb{E}\int_{Q}Cz\,Bz-2\mathbb{E}\int_{Q}|M_{h}z|^{2}\,dt.$
The next step is to provide an estimate for the right-hand side of (B.4).
Remark B.1.
The decomposition of (B.2) differs from that presented in EstimateCarlemanStochasticControl , not only due to the sign change, but also because the corrector, which is typically added when the dimension is greater than one, is included as an additional term rather than as a component of $M_{h}z$. As a result, the estimate for the cross-product includes an additional term that must be considered. Moreover, in the estimate (B.4), Young’s inequality is applied in a way that preserves the term $|C(z)|^{2}$, which will play a crucial role in the final estimate of the Carleman inequality. These are some of the reasons for the study of this new configuration in order to obtain the Carleman inequality.
B.2. An estimate for the cross-product
@@
(B.5) $2\mathbb{E}\int_{Q}CzBz=2\mathbb{E}\sum_{i=1}^{6}\sum_{j=1}^{3}\int_{Q}C_{i}zB_{j}z\triangleq\,\sum_{i=1}^{6}\sum_{j=1}^{3}I_{ij}.$
The procedure for estimating the product (B.5) is analogous to that in EstimateCarlemanStochasticControl , except for a change of sign and consideration of Remark B.1. Consequently, we divide the estimate into three groups: terms involving the differential $dz$, products of the additional terms, and terms involving the differential $dt$.
We begin with the terms involving the differential $dt$. Specifically, consider the terms $I_{i}$ for $i=1,2,3,\dots,5$, whose estimates can be found in EstimateCarlemanStochasticControl , up to a sign change. For the additional term $I_{6}$, which does not appear in EstimateCarlemanStochasticControl , its estimate is provided in C. Consequently, we obtain the following preliminary estimate:
Lemma B.2.
@@
$\begin{split}X_{1}\triangleq\,&\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q_{i}^{\ast}}T\theta(t)\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right)\\ &+\mathbb{E}\int_{Q}(\mathcal{O}_{\lambda}((sh)^{2})+T\mathcal{O}_{\lambda}(1)+s\lambda\varphi\mathcal{O}(1))\,|dz|^{2}\\ &+\mathbb{E}\int_{Q}T(\theta\mathcal{O}_{\lambda}((sh)^{2})-2s\varphi(\lambda^{3}\beta|\nabla_{\gamma}\psi|^{2}+\lambda^{2}\mathcal{O}(1)))+\mathcal{O}_{\lambda}(1))\,|z|^{2}dt\end{split}$
@@
$\begin{split}Y_{1}\triangleq\,&-\left.\mathbb{E}\int_{\mathcal{M}}[s^{2}\mathcal{O}_{\lambda}(1)+2s(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))]|z|^{2}\right|_{0}^{T}+\sum_{i=1}^{n}\left.\mathbb{E}\int_{\mathcal{M}_{i}}\gamma_{i}|D_{i}z|^{2}\right|_{0}^{T}\\ &\sum_{i=1}^{n}\mathbb{E}\left.\int_{\mathcal{M}_{i}^{*}}\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\right|_{0}^{T}+\mathbb{E}\left.\int_{\mathcal{M}}(\mathcal{O}_{\lambda}((sh)^{2})+T\theta^{2}\mathcal{O}(1))\,|z|^{2}\right|_{0}^{T}.\end{split}$
Now, let us consider the products of the additional terms. In this case, the new terms $I_{62}$ and $I_{63}$ appear, whose estimates are provided in C. For the remaining terms, the estimates remain unchanged compared to EstimateCarlemanStochasticControl , since the weight functions still satisfy the same properties, namely $|\theta(t)|\leq C$ and $|\theta_{t}(t)|\leq C$ for some constant $C>0$ and all $t\in[0,T]$. Therefore, we obtain the following estimate:
Lemma B.3.
@@
$X_{3}\triangleq\,\mathbb{E}\int_{Q}s|\mathcal{O}_{\lambda}((sh)^{2})|\,|z|^{2}\,dt-E\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s|\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}z|^{2}\,dt$
@@
Finally, the terms $I_{12}$, $I_{22}$ and $I_{32}$ behave analogously to those in the deterministic setting studied in boyer-2014 , as the time variable does not play a significant role. For this reason, we omit a detailed proof of the following lemma, since the only difference with respect to boyer-2014 lies in our notation.
Lemma B.4.
@@
$X_{2}\triangleq\,\sum_{i=1}^{n}\mathbb{E}\int_{Q}|s\lambda\varphi\mathcal{O}(1)+\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}A_{i}z|^{2}\,dt$
@@
$Y_{2}\triangleq\,$ $\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}\left(-2s\lambda\varphi(\gamma_{i})^{2}\partial_{i}\psi+s\mathcal{O}_{\lambda}(sh)+h\mathcal{O}_{\lambda}(sh)\right)t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt+$
@@
Then by Lemma B.2-B.4 we obtain the following
(B.6) $\begin{split}2\mathbb{E}&\int_{Q}CzBz\geq\,\mathbb{E}\int_{Q}2s^{3}\lambda^{4}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}2s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}\,|D_{i}z|^{2}\,dt\\ &-\sum_{i=1}^{n}\mathbb{E}\int_{Q^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}-\sum_{j=1}^{3}X_{j}-\sum_{j=1}^{2}Y_{j}.\end{split}$
To give an estimate of the right-hand side of (B.4), we need the following estimate $M_{h}z$. The proof can be adapted from Lemma 4.2 in boyer-2010-1d-elliptic and the estimation of $\Phi$ in zhao:2024 .
Lemma B.5.
@@
In addition, we will need the following lemma, where its proof proceeds analogously to the proof of Lemma 3.11 in boyer-2014 .
Lemma B.6.
@@
$\begin{split}&X_{3}\triangleq\,\sum_{i\neq=1}^{n}\left(\mathbb{E}\int_{Q_{ij}^{\ast}}h\mathcal{O}_{\lambda}(sh)\,|D^{2}_{ij}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}h\mathcal{O}(sh)|D_{i}z|^{2}\,dt\right)\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|D^{2}_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}h\mathcal{O}_{\lambda}(sh)|D_{i}z|^{2}\,dt\right)\end{split}$
@@
$Y_{3}\triangleq\,\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}h\mathcal{O}(sh)t_{r}^{i}(|D_{i}z|^{2})\,dt.$
Combining the lemma B.5 and B.6 with (B.6), we see that for $sh\leq\varepsilon$, there exist $\lambda_{1}\geq 1$, and $\varepsilon_{1}(\lambda)>0$ such that for $\lambda\geq\lambda_{1}$ and $0<sh\leq\varepsilon_{1}(\lambda)$, we have
@@
(B.8) $\begin{split}\tilde{X}\triangleq&\sum_{l=1}^{2}X_{l}+Y_{l}-\mathbb{E}\int_{Q}s^{3}\lambda^{3}\varphi^{3}\mathcal{O}(1)|z|^{2}\,dt\end{split}-\mathbb{E}\int_{Q}s^{2}\lambda^{4}\varphi^{2}\mathcal{O}(1)\,|z|^{2}\,dt$
If we fix $\lambda=\lambda_{1}$, we can choose $\varepsilon_{0}$ and $h_{0}$ to be sufficiently small, with $0<\varepsilon\leq\varepsilon(\lambda_{1})$, and $s_{0}\geq 1$ sufficiently large, so that for $s\geq s_{0}$, $0<h\leq h_{0}$, and $sh\leq\varepsilon_{0}$, we obtain
@@
(B.10) $\begin{split}\tilde{Y}\triangleq\,\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}2s\lambda_{1}\varphi(\gamma_{i})^{2}\partial_{i}\psi\,&t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt\\ &+C_{s_{0},\epsilon}\left(\left.-\sum_{i=1}^{n}\mathbb{E}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}z|^{2}\right|_{0}^{T}+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,|z|^{2}\right|_{0}^{T}\right)\end{split}$
@@
$\tilde{Y}=\,C_{s_{0},\epsilon}\left(\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}s\lambda t_{r}^{i}(|D_{i}z|^{2})\,dt-\left.\mathbb{E}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}z|^{2}\right|_{t=T}+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}\right).$
Therefore, we write (B.9) as
(B.11) $\begin{split}&\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda_{1}^{2}\varphi^{2}+s\lambda_{1}^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}+\mathbb{E}\int_{Q}|Cz|^{2}\,dt\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda^{2}_{1}|\nabla_{\gamma}\psi|^{2}\varphi\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi|\nabla_{\gamma}\psi|^{2}|A_{i}D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}\right)\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}s\lambda_{1}t_{r}^{i}(|D_{i}z|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}\right).\end{split}$
B.3. Returning to the original variable
We now return to our original variable. The methodology in this section departs from the approach in EstimateCarlemanStochasticControl . Therefore, as a first step, we establish the following result for the terms on the left-hand side of the desired Carleman estimate, which corresponds to the left-hand side of inequality (B.11).
Lemma B.7.

### v2

@@ A.1. Conjugated operator
For all i=1,,n.i{=1,...,n.}, let us consider the functions $\gamma_{i}$ such that $\mbox{reg}(\gamma)<\mbox{reg}^{0}$ and the following notation
$\nabla_{\gamma}f:=(\sqrt{\gamma_{1}}D_{1}f_{1},\cdots,\sqrt{\gamma_{n}}D_{n}f_{n})\quad\text{and}\,\quad\Delta_{\gamma}f:=\sum_{i=1}^{n}\gamma_{i}\partial_{x_{i}}^{2}f.$
Let $\mathcal{P}(w):=\,dw-\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}w\right)dt=fdt+gdB(t)$. With respect to this weight function, the first step is to consider the change of variable $w=\rho z$. Our first task is to split the conjugate operator
@@
into simple terms that we will estimate separately. Using Itô’s formula (cf. (Klebaner, 2012, Theorem 4.18)) and notice that $r\rho=1$, we obtain the following result
@@
(A.1) $\begin{split}rd(\rho z)-r\sum_{i=1}^{n}D_{i}\left(\gamma_{i}D_{i}(\rho z)\right)dt+R_{h}z\,dt=&B_{1}z+B_{2}z\,dt+C_{1}z\,dt+C_{2}z\,dt+C_{3}z\,dt,\end{split}$
@@
$\begin{split}R_{h}z:=\,&\sum_{i=1}^{n}\left(h\mathcal{O}(1)rD_{i}^{2}\rho+D_{i}\gamma_{i}rA_{i}D_{i}\rho\right)A_{i}^{2}z+\sum_{i=1}^{n}h_{i}\mathcal{O}(1)rD_{i}A_{i}\rho D_{i}A_{i}z\\ &+\sum_{i=1}^{n}\frac{h^{2}}{4}D_{i}\gamma_{i}rD_{i}^{2}\rho D_{i}A_{i}z+\sum_{i=1}^{n}\frac{h^{2}}{4}D_{i}\gamma_{i}\,rD_{i}A_{i}\rho D_{i}^{2}z.\end{split}$
@@
$C_{4}z:=\,-\frac{h^{2}}{4}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}(rD_{i}^{2}\rho)A_{i}z),\quad C_{5}z:=\,-\frac{h^{2}}{4}\sum_{i=1}^{n}D_{i}(D_{i}(\gamma_{i}rD^{2}_{i}\rho)A_{i}z)$
$C_{6}z=-B_{3}z:=\,-2s\Delta_{\gamma}\varphi\,z$
in (A.1) we have the following identity in $Q$
(A.2) $r\mathcal{P}(\rho z)=Cz\,dt+Bz-M_{h}z\,dt,$
@@
Thus, our next task is to estimate the cross-product $2\mathbb{E}\int_{Q}CzBz=2\mathbb{E}\sum_{i=1}^{6}\sum_{j=1}^{3}\int_{Q}C_{i}zB_{j}z$ and to bound the left-hand side of (A.2) with respect to the term $M_{h}z$. Indeed, We notice that from (A.2), it holds that
(A.3) $2r\mathcal{P}(q)\,Cz=2|Cz|^{2}dt+2Cz\,Bz-2M_{h}z\,Cz\,dt.$
Hence, combining the definition of $r\mathcal{P}(q)$ with (A.3), it follows that
@@
(A.4) $2\mathbb{E}\int_{0}^{T}|rf|^{2}\,dt\geq\mathbb{E}\int_{Q}|Cz|^{2}dt+2\mathbb{E}\int_{Q}Cz\,Bz-2\mathbb{E}\int_{Q}|M_{h}z|^{2}\,dt.$
The next step is to provide an estimate for the right-hand side of (A.4).
Remark A.1.
The decomposition of (A.2) differs from that presented in Lecaros et al. (2026), not only due to the sign change, but also because the corrector, which is typically added when the dimension is greater than one, is included as an additional term rather than as a component of $M_{h}z$. As a result, the estimate for the cross-product includes an additional term that must be considered. Moreover, in the estimate (A.4), Young’s inequality is applied in a way that preserves the term $|C(z)|^{2}$, which will play a crucial role in the final estimate of the Carleman inequality. These are some of the reasons for the study of this new configuration in order to obtain the Carleman inequality.
A.2. An estimate for the cross-product
@@
(A.5) $2\mathbb{E}\int_{Q}CzBz=2\mathbb{E}\sum_{i=1}^{6}\sum_{j=1}^{3}\int_{Q}C_{i}zB_{j}z:=\,\sum_{i=1}^{6}\sum_{j=1}^{3}I_{ij}.$
The procedure for estimating the product (A.5) is analogous to that in Lecaros et al. (2026), except for a change of sign and consideration of Remark A.1. Consequently, we divide the estimate into three groups: terms involving the differential $dz$, products of the additional terms, and terms involving the differential $dt$.
We begin with the terms involving the differential $dt$. Specifically, consider the terms $I_{i}$ for $i=1,2,3,\dots,5$, whose estimates can be found in Lecaros et al. (2026), up to a sign change. For the additional term $I_{6}$, which does not appear in Lecaros et al. (2026), its estimate is provided in Appendix B. Consequently, we obtain the following preliminary estimate:
Lemma A.2.
@@
$\begin{split}X_{1}:=\,&\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q_{i}^{\ast}}T\theta(t)\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right)\\ &+\mathbb{E}\int_{Q}(\mathcal{O}_{\lambda}((sh)^{2})+T\mathcal{O}_{\lambda}(1)+s\lambda\varphi\mathcal{O}(1))\,|dz|^{2}\\ &+\mathbb{E}\int_{Q}T(\theta\mathcal{O}_{\lambda}((sh)^{2})-2s\varphi(\lambda^{3}\beta|\nabla_{\gamma}\psi|^{2}+\lambda^{2}\mathcal{O}(1)))+\mathcal{O}_{\lambda}(1))\,|z|^{2}dt\end{split}$
@@
$\begin{split}Y_{1}:=\,&-\left.\mathbb{E}\int_{\mathcal{M}}[s^{2}\mathcal{O}_{\lambda}(1)+2s(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))]|z|^{2}\right|_{0}^{T}+\sum_{i=1}^{n}\left.\mathbb{E}\int_{\mathcal{M}_{i}}\gamma_{i}|D_{i}z|^{2}\right|_{0}^{T}\\ &\sum_{i=1}^{n}\mathbb{E}\left.\int_{\mathcal{M}_{i}^{*}}\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\right|_{0}^{T}+\mathbb{E}\left.\int_{\mathcal{M}}(\mathcal{O}_{\lambda}((sh)^{2})+T\theta^{2}\mathcal{O}(1))\,|z|^{2}\right|_{0}^{T}.\end{split}$
Now, let us consider the products of the additional terms. In this case, the new terms $I_{62}$ and $I_{63}$ appear, whose estimates are provided in Appendix B. For the remaining terms, the estimates remain unchanged compared to Lecaros et al. (2026), since the weight functions still satisfy the same properties, namely $|\theta(t)|\leq C$ and $|\theta_{t}(t)|\leq C$ for some constant $C>0$ and all $t\in[0,T]$. Therefore, we obtain the following estimate:
Lemma A.3.
@@
$X_{3}:=\,\mathbb{E}\int_{Q}s|\mathcal{O}_{\lambda}((sh)^{2})|\,|z|^{2}\,dt-E\int_{Q}s\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s|\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}z|^{2}\,dt$
@@
Finally, the terms $I_{12}$, $I_{22}$ and $I_{32}$ behave analogously to those in the deterministic setting studied in Boyer and Le Rousseau (2014), as the time variable does not play a significant role. For this reason, we omit a detailed proof of the following lemma, since the only difference with respect to Boyer and Le Rousseau (2014) lies in our notation.
Lemma A.4.
@@
$X_{2}:=\,\sum_{i=1}^{n}\mathbb{E}\int_{Q}|s\lambda\varphi\mathcal{O}(1)+\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}(sh)+s\mathcal{O}_{\lambda}((sh)^{2})|\,|D_{i}A_{i}z|^{2}\,dt$
@@
$Y_{2}:=\,$ $\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}\left(-2s\lambda\varphi(\gamma_{i})^{2}\partial_{i}\psi+s\mathcal{O}_{\lambda}(sh)+h\mathcal{O}_{\lambda}(sh)\right)t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt+$
@@
Then by Lemma A.2-A.4 we obtain the following
(A.6) $\begin{split}2\mathbb{E}&\int_{Q}CzBz\geq\,\mathbb{E}\int_{Q}2s^{3}\lambda^{4}\varphi^{3}|\nabla_{\gamma}\psi|^{4}\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}2s\lambda^{2}\varphi\gamma_{i}|\nabla_{\gamma}\psi|^{2}\,|D_{i}z|^{2}\,dt\\ &-\sum_{i=1}^{n}\mathbb{E}\int_{Q^{\ast}}\gamma_{i}|D_{i}(dz)|^{2}+\mathbb{E}\int_{Q}(s^{2}\lambda^{2}\varphi^{2}+s\lambda^{2}\varphi)|\nabla_{\gamma}\psi|^{2}|dz|^{2}-\sum_{j=1}^{3}X_{j}-\sum_{j=1}^{2}Y_{j}.\end{split}$
To give an estimate of the right-hand side of (A.4), we need the following estimate $M_{h}z$. The proof can be adapted from Lemma 4.2 in Boyer et al. (2010a) and the estimation of $\Phi$ in Zhao (2025).
Lemma A.5.
@@
In addition, we will need the following lemma, where its proof proceeds analogously to the proof of Lemma 3.11 in Boyer and Le Rousseau (2014).
Lemma A.6.
@@
$\begin{split}&X_{3}:=\,\sum_{i\neq=1}^{n}\left(\mathbb{E}\int_{Q_{ij}^{\ast}}h\mathcal{O}_{\lambda}(sh)\,|D^{2}_{ij}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}h\mathcal{O}(sh)|D_{i}z|^{2}\,dt\right)\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|A_{i}D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}h\mathcal{O}_{\lambda}(sh)|D^{2}_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}h\mathcal{O}_{\lambda}(sh)|D_{i}z|^{2}\,dt\right)\end{split}$
@@
$Y_{3}:=\,\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}h\mathcal{O}(sh)t_{r}^{i}(|D_{i}z|^{2})\,dt.$
Combining the lemma A.5 and A.6 with (A.6), we see that for $sh\leq\varepsilon$, there exist $\lambda_{1}\geq 1$, and $\varepsilon_{1}(\lambda)>0$ such that for $\lambda\geq\lambda_{1}$ and $0<sh\leq\varepsilon_{1}(\lambda)$, we have
@@
(A.8) $\begin{split}\tilde{X}:=&\sum_{l=1}^{2}X_{l}+Y_{l}-\mathbb{E}\int_{Q}s^{3}\lambda^{3}\varphi^{3}\mathcal{O}(1)|z|^{2}\,dt\end{split}-\mathbb{E}\int_{Q}s^{2}\lambda^{4}\varphi^{2}\mathcal{O}(1)\,|z|^{2}\,dt$
If we fix $\lambda=\lambda_{1}$, we can choose $\varepsilon_{0}$ and $h_{0}$ to be sufficiently small, with $0<\varepsilon\leq\varepsilon(\lambda_{1})$, and $s_{0}\geq 1$ sufficiently large, so that for $s\geq\tau_{0}$, $0<h\leq h_{0}$, and $sh\leq\varepsilon_{0}$, we obtain
@@
(A.10) $\begin{split}\tilde{Y}:=\,\sum_{i=1}^{n}\mathbb{E}\int_{\partial_{i}Q}2s\lambda_{1}\varphi(\gamma_{i})^{2}\partial_{i}\psi\,&t_{r}^{i}(|D_{i}z|^{2})\nu_{i}\,dt\\ &+\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\left.-\sum_{i=1}^{n}\mathbb{E}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}z|^{2}\right|_{0}^{T}+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,|z|^{2}\right|_{0}^{T}\right)\end{split}$
@@
$\tilde{Y}\leq\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{\mathcal{M}}s^{2}|z|\bigg|_{t=T}\right).$
Therefore, we write (A.9) as
(A.11) $\begin{split}\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}&\varphi^{3}|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda_{1}^{2}\varphi^{2}+s\lambda_{1}^{2}\varphi)|dz|^{2}+\mathbb{E}\int_{Q}|Cz|^{2}\,dt\\ &+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda^{2}_{1}\varphi\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi|A_{i}D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}|D_{i}(dz)|^{2}\right)\\ &\leq\,\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}}s^{3}\lambda^{4}_{1}\varphi^{3}\,|z|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}_{i}^{\ast}}s\lambda^{2}_{1}\varphi\,|D_{i}z|^{2}\,dt\\ &+\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}\right).\end{split}$
A.3. Returning to the original variable
We now return to our original variable. The methodology in this section departs from the approach in Lecaros et al. (2026). Therefore, as a first step, we establish the following result for the terms on the left-hand side of the desired Carleman estimate, which corresponds to the left-hand side of inequality (A.11).
Lemma A.7.